im having trouble with algebra word problems im stuck... A total of $800 was split into wo investments. Part paid 9%, and the remainder paid 6%. If the annual interest from the two investments was $67.50, how much was invested at each rate?
Answer by jorel1380(3719)

p+q=800
9p+6q=6750
6p+6q=4800
3p=1950
p=650
q=150
$650 was invested at 9%, $150 was invested at 6%..